# CSEval: Towards Automated, Multi-Dimensional, and Reference-Free Counterspeech Evaluation using Auto-Calibrated LLMs

### v1

@@ Abstract
Counterspeech has been popularly shown as an effective approach to counter online hate speech, leading to increasing research interest in automated counterspeech generation using language models. However, this field lacks standardised evaluation protocols and robust automated evaluation metrics that align with human judgement. Current automatic evaluation methods, primarily based on similarity metrics, do not effectively capture the complex and independent attributes of counterspeech quality, such as contextual relevance, aggressiveness, or argumentative coherence. This has led to an increased dependency on labor-intensive human evaluations to assess automated counter-speech generation methods. To address these challenges, we introduce CSEval, a novel dataset and framework for evaluating counterspeech quality across four dimensions: contextual-relevance, aggressiveness, argument-coherence, and suitableness. Furthermore, we propose Auto-Calibrated COT for Counterspeech Evaluation (ACE), a prompt-based method with auto-calibrated chain-of-thoughts (CoT) for scoring counterspeech using large language models. Our experiments show that ACE outperforms traditional metrics like ROUGE, METEOR, and BertScore in correlating with human judgement, indicating a significant advancement in automated counterspeech evaluation. 111 Warning: The content in this paper may be upsetting or offensive.
@@ 1 Introduction
The emergent capabilities of LLMs such as instruction tuning Wei et al. (2022), and preference tuning through human feedback Ouyang et al. (2022), and as Chain-of-Thought (CoT) reasoning Wei et al. (2023) have presented a promising direction towards NLG evaluation based on LLMs. Recent studies have proposed to use LLMs as reference-free NLG evaluators directly, offering a better human-aligned evaluation compared to traditional methods. Wang et al. (2023); Fu et al. (2023); Liu et al. (2023a). However, the validity and reliability of using LLMs as NLG evaluators have yet to be systematically investigated in automated counterspeech generation Chung et al. (2024). Specifically, there is an issue with how automated metrics are evaluated themselves. Currently, there is no dataset with human judgments evaluating specific quality aspects of generated counterspeech.
We address these gaps in complementary ways. First, we build CSEval, the largest and most diverse(in terms of model types) collection of human judgments of model-generated counterspeech across four quality aspects: contextual-relevance, aggressiveness, argument-coherence, and suitableness. We release aligned model outputs for five of the current state-of-the-art (SoTA) models on counterspeech generation. We benchmark CSEval on multiple reference and aspect-based automated evaluation methods. Finally, we propose Auto-Calibrated COT for Counterspeech Evaluation (ACE), a prompt-based method with auto-calibrated CoT to score each quality dimension of the counterspeech. Experimental results show that our proposed approach outperforms the traditional and LLM-based evaluation methods in terms of correlation with human judgment.
To summarize our contribution, in this paper, we – (i) introduce CSEval, the largest and most diverse collection of human judgments of model-generated counterspeech across four quality aspects: context relevance, aggressiveness, argument coherence, and suitableness; (ii) benchmark the CSEval dataset across multiple popular automatic evaluation methods and find that none of them reliably measure the quality of counterspeech, showing poor correlation with human-judgments; (iii) propose ACE, a prompt-based method with auto-calibrated CoTs to score each quality dimension of the counterspeech, and show that it outperforms the traditional and LLM-based evaluation strategies which are correlated with the human judgment  [^2] The source code and datasets are available at https://github.com/AmeyHengle/cs-eval.
@@ 2 Related Work
With the recent progress in this domain, automated evaluation of model-generated counterspeech remains an open problem. Generally speaking, almost all of the previous studies Zhu and Bhat (2021); Gupta et al. (2023); Saha et al. (2022); Hengle et al. (2024) extensively employ lexical overlap Papineni et al. (2002b); Lin (2004); Banerjee and Lavie (2005), semantic similarity Zhang et al. (2020a), and diversity Li et al. (2016) metrics for counterspeech evaluation. However, reference-based metrics have been consistently shown to have poor correlation with human judgments, especially for evaluating the quality of open-ended generation tasks Zhong et al. (2022); Fu et al. (2023); Li et al. (2024). Furthermore, these metrics are incapable of representing the aspect-level quality of a counterspeech, such as soundness, relevance, or non-toxicity Chung et al. (2024). Due to these reasons, research in counterspeech generation is heavily dependent on human evaluation, which is often very costly to conduct and difficult to scale.
With their growing reasoning and generation capabilities, LLMs are now being considered as an alternative to human evaluation, showing a high degree of correlation with human judgement on evaluating tasks like summarisation, dialog-generation, and code generation Wang et al. (2023); Liu et al. (2023b); Fu et al. (2023); Liu et al. (2023a); Li et al. (2024). Fu et al. (2023) propose GPTScore, an automated framework that evaluates text quality with pre-trained generative models like GPT3. Kocmi and Federmann (2023) propose to use GPT-based models for scoring machine translation outputs. Li et al. (2024) conduct an extensive survey on the usage of LLMs for NLG evaluation. Recently, Liu et al. (2023a) improved correlation with human judgement by generating detailed evaluation steps using the CoT reasoning capability of GPT-4. Our proposed approach, ACE uses a similar strategy. However, instead of relying on model-generated CoT, we actively calibrate the LLM based on correlation with human scores.
@@ 3 Dataset
We build and release CSEval, a benchmark dataset for reference-free and multi-dimensional counterspeech evaluation. CSEval contains expert human assessments of $7,926$ model-generated CS, across four quality dimensions (or aspects). This section provides a detailed overview of the data curation and annotation process, evaluation dimensions, and the models used in CSEval.
@@ 3.3 Models
We include counterspeeches generated using both supervised and prompt-based methods. In terms of supervised methods, we include three popular counterspeech generation models – Generate-Prune-Select (GPS) Zhu and Bhat (2021), DialoGPT Zhang et al. (2020b), and QUARC Gupta et al. (2023). To further broaden our scope, we also consider prompting baselines leveraging the capabilities of LLMs. Specifically, we include counterspeeches generated by GPT-3.5-Turbo (ChatGPT) and GPT-4 Ouyang et al. (2022), prompted in both zero- and few-shot fashion respectively. Further details surrounding model training and prompting are provided under Appendix A.1.1.
@@ 3.4 Annotation Process
We used Krippendorff’s alpha coefficient Klaus (2011) to measure the inter-annotator agreement of the expert annotations. For the first round of annotations, we obtained a decent inter-annotator interval of Krippendorff’s (averaged across the four dimensions) of $0.473$. However, the second round of annotations increases the inter-annotator agreement, achieving an average Krippendoff coefficient of $0.681$. For more insights, we calculate the standard deviations of annotator scores within the respective groups and display histograms of those statistics in Appendix Figure 3. The histograms show that while the second round of annotations has reduced the standard deviation of scores and significantly enhanced inter-annotator agreement, suitableness remains the most contentious dimension among experts. Further details about the annotation process are provided under Appendix Section A.
Figure 2: An overview of the multi-phase auto-calibration framework of ACE, including the generation and refinement of CoT instructions, evaluation criteria formulation, and the iterative calibration process aligned with expert human judgement.
@@ 4 Proposed Method
Auto Evaluation Steps (CoT): CoT is a prompt engineering technique that encourages LLMs to break down a complex task into a sequence of intermediate steps and has been shown to improve LLMs in their reasoning abilities and generate more accurate and informative responses for multi-step problems, such as arithmetic, common-sense, and symbolic reasoning Wang et al. (2023); Zheng et al. (2023). For evaluation tasks, some CoTs need more detailed instructions beyond the simple definition, and it is time-consuming to manually design such evaluation steps for each task. Liu et al. (2023a) found that an LLM can generate such evaluation steps by itself, and proposed an auto-CoT approach to evaluate NLG tasks. However, this approach may result in sub-optimal and misaligned CoT where the scoring guidelines are absent and only output categorical ranges (e.g., 0-5) are provided. LLM-based evaluators are shown to suffer from insufficient prompting, resulting in inconsistent and misaligned evaluations Lu et al. (2023). Furthermore, it is non-trivial to define the CoT instructions (or evaluation steps) manually, as it is time-consuming and requires expert-level domain knowledge to assign rubrics and prevent personal bias. Recently, Liu et al. (2023b) proposed a framework to auto-calibrate LLM prompts based on a validation set of human judgments. Drawing inspiration from this, we propose a data-driven approach to calibrating the most optimal CoT instructions (or evaluation steps) for a given dimension. As illustrated in Figure 2, our method follows a two-phase procedure.
In the first phase, we begin by constructing a validation set $D^{*}$ of $500$ model outputs by randomly sampling from the CSEval dev set, denoted by $D$. After this, we utilize the following instructions and the in-context learning capability of LLMs to independently infer scoring CoT $C$ from few-shot exemplars. One crucial part here is to ensure the diversity of recalled CoT. To mitigate the label bias and position bias of in-context learning Liu et al. (2023b), we construct Monte-Carlo samples from $D^{*}$ to sample few-shot in-context examples.
Given an arbitrary sample $d_{i}\in D$, the prompt template for Chain of Thought (CoT) drafting $T_{D}$, evaluation aspect $a$ (e.g., relevance, coherence), and a model $LLM(\cdot)$, the quality dimension of $d_{i}$ is evaluated as $\hat{s}_{i,a}=LLM(T_{D}(d_{i},C,a))$. With this setup in place, a corresponding candidate CoT can be inferred as:
@@
where $D_{s}=\bigcup_{i}(d_{i}^{*},s_{i,a})\subset D^{*}$ denotes the few-shot exemplar subset, and $P_{\theta}$ represents the probability distribution modeled by the parameters $\theta$ of the LLM.
To allow for diversity in candidate CoT prompts, we apply temperature sampling to draw scoring CoT from the LLM. The temperature values for temperature sampling are dynamically selected from a range of 00 to $0.5$, for generation diversity while curating the candidate CoT. Example prompt templates are provided in Appendix B. Following this procedure, we obtain the initial set of candidate CoTs for evaluation and refinement.
In the second phase, we iteratively refine the candidate CoT pool obtained from the previous phase using $D^{*}$. Inferred from various few-shot examples, CoTs within the initial draft sets may be sub-optimal or contain potential bias (e.g., to particular exemplar labels). In order to filter out high-quality candidates, we first revisit them using $D^{*}$, and then select the top-performing candidates according to their human correlation. To mitigate disagreements between human experts and the drafted CoT, we prompt LLMs to refine the previously generated CoT by providing them samples with strong disagreement in their correlation scores. That is, we identify the top-3 instances with the highest absolute difference between the predicted and human scores and use these as few-shot examples for the CoT refinement process.. This ensures that the candidate CoTs are aligned more closely with human ratings and reduces potential biases in the final prompts.
@@
    Modification: Adjusting some parts of CoT to increase its correlation.
@@
    Paraphrase: If a CoT is good enough, paraphrase it to make it more clear and more concise.
@@
    Adding aspects or details: When the LLM finds new scoring rules that are not present in the current CoT, consider adding them as a new line to the current CoT but make sure not to make the CoT too long and redundant.
@@
As illustrated in Figure 2, after obtaining refined candidate CoT, we first filter them with $D^{*}$ and then combine them with the pre-filtered draft CoT to obtain a final calibrated set of scoring rules. The detailed pseudo-code for CoT selection and refinement is provided in Algorithm 1.
@@
GPT-4 (ACE) $0.687$ $0.637$ $0.447$ $0.425$ $0.600$ $0.506$ $0.567$ $0.504$ $0.575$ $0.518$
$\Delta_{\mathrm{GPT-4(ACE)(Ours)}-\mathrm{BestMethod}}$ $\uparrow 0.062$ $\uparrow 0.065$ $\uparrow 0.068$ $\uparrow 0.075$ $\uparrow 0.015$ $\uparrow 0.019$ $\uparrow 0.018$ $\uparrow 0.040$ $\uparrow 0.048$ $\uparrow 0.049$
Table 1: Sample-level Spearman ($\rho$) and Kendall-Tau ($\tau$) correlations of different metrics on CSEval benchmark. For any given evaluation aspect (Relevance, for example), the correlation scores are computed against their respective human ratings. LLM-based evaluators are reported in three settings: zero-shot, chain-of-thoughts (G-EVAL), and auto-calibrated chain-of-thoughts (ACE). Automatic and human-evaluation scores were normalised to a scale of (0,1) before computing correlation.

### v2

@@ Abstract
Counterspeech has emerged as a popular and effective strategy for combating online hate speech, sparking growing research interest in automating its generation using language models. However, the field still lacks standardised evaluation protocols and reliable automated evaluation metrics that align with human judgement. Current automatic evaluation methods, primarily based on similarity metrics, do not effectively capture the complex and independent attributes of counterspeech quality, such as contextual relevance, aggressiveness, or argumentative coherence. This has led to an increased dependency on labor-intensive human evaluations to assess automated counter-speech generation methods. To address these challenges, we introduce CSEval, a novel dataset and framework for evaluating counterspeech quality across four dimensions: contextual-relevance, aggressiveness, argument-coherence, and suitableness. Furthermore, we propose Auto-Calibrated COT for Counterspeech Evaluation (Auto-CSEval), a prompt-based method with auto-calibrated chain-of-thoughts (CoT) for scoring counterspeech using large language models. Our experiments show that Auto-CSEval outperforms traditional metrics like ROUGE, METEOR, and BertScore in correlating with human judgement, indicating a significant improvement in automated counterspeech evaluation. 111 Warning: The content in this paper may be upsetting or offensive.
@@ 1 Introduction
The emergent capabilities of LLMs such as instruction tuning Wei et al. (2022), and preference tuning through human feedback Ouyang et al. (2022), and as Chain-of-Thought (CoT) reasoning Wei et al. (2023) have presented a promising direction towards NLG evaluation based on LLMs. Recent studies have proposed using LLMs as reference-free NLG evaluators directly, offering a better human-aligned evaluation compared to traditional methods. Wang et al. (2023); Fu et al. (2023); Liu et al. (2023a); Jones et al. (2024). However, the validity and reliability of using LLMs as NLG evaluators have yet to be systematically investigated in automated counterspeech generation Chung et al. (2024). Specifically, there is an issue with how automated metrics are evaluated themselves. Jones et al. (2024) took an initial step in this direction by proposing a dataset and framework for reference-free counterspeech evaluation. However, they lack expert annotations and have a relatively small test size (180 samples) Jones et al. (2024). Thus, there is no dataset with expert human judgements evaluating specific quality aspects of model-generated counterspeech.
We address these gaps in complementary ways. First, we build CSEval, a large and diverse collection (in terms of model-types) of human judgements of model-generated counterspeech across four quality aspects: contextual-relevance, aggressiveness, argument-coherence, and suitableness. We release aligned model outputs for five of the current state-of-the-art (SoTA) models on counterspeech generation. We benchmark CSEval on multiple reference-based as well as reference-free evaluation methods. Finally, we propose Auto Calibrated Chain-of-Thoughts for Counterspeech Evaluation (Auto-CSEval), a prompt-based method with auto-calibrated evaluation steps (CoT) to score each quality dimension of the counterspeech. Experimental results show that our proposed approach outperforms the traditional and LLM-based evaluation methods in terms of correlation with human judgement.
To summarize our contribution, in this paper, we – (i) introduce CSEval, the largest and most diverse collection of human judgments of model-generated counterspeech across four quality aspects: context relevance, aggressiveness, argument coherence, and suitableness; (ii) benchmark the CSEval dataset across multiple popular automatic evaluation methods and find that none of them reliably measure the quality of counterspeech, showing poor correlation with human-judgments; (iii) propose Auto-CSEval, a prompt-based method with auto-calibrated CoTs to score each quality dimension of the counterspeech, and show that it outperforms the traditional and LLM-based evaluation strategies which are correlated with the human judgment  [^2] The source code and datasets are available at https://github.com/AmeyHengle/cs-eval.
@@ 2 Related Work
With the recent progress in this domain, automated evaluation of model-generated counterspeech remains an open problem. Generally speaking, almost all of the previous studies Zhu and Bhat (2021); Gupta et al. (2023); Saha et al. (2022); Hengle et al. (2024) extensively employ lexical overlap Papineni et al. (2002b); Lin (2004); Banerjee and Lavie (2005), semantic similarity Zhang et al. (2020a), and diversity Li et al. (2016) metrics for counterspeech evaluation. However, reference-based metrics have been consistently shown to have poor correlation with human judgements, especially for evaluating the quality of open-ended generation tasks Zhong et al. (2022); Fu et al. (2023); Li et al. (2024). Furthermore, these metrics are incapable of representing the aspect-level quality of a counterspeech, such as soundness, relevance, or non-toxicity Chung et al. (2024). Due to these reasons, research in counterspeech generation is heavily dependent on human evaluation, which is often very costly to conduct and difficult to scale.
With their growing reasoning and generation capabilities, LLMs are now being considered as an alternative to human evaluation, showing a high degree of correlation with human judgement on evaluating tasks like summarisation, dialog-generation, and code generation Wang et al. (2023); Liu et al. (2023b); Fu et al. (2023); Liu et al. (2023a); Li et al. (2024). Fu et al. (2023) propose GPTScore, an automated framework that evaluates text quality with pre-trained generative models like GPT3. Kocmi and Federmann (2023) propose to use GPT-based models for scoring machine translation outputs. Li et al. (2024) conduct an extensive survey on the usage of LLMs for NLG evaluation. Recently, Jones et al. (2024) introduced a reference-free evaluation framework that leverages LLMs to assess counterspeech quality across multiple dimensions. They showed that LLM-based evaluation aligns more closely with human judgement compared to traditional methods. Liu et al. (2023a) propose a prompt-based method to further improve correlation with human judgement by generating detailed evaluation steps using the CoT reasoning capability of GPT-4. Our proposed approach, Auto-CSEval, builds on both these works. However, instead of relying solely on model-generated CoT reasoning like Liu et al. (2023a), we actively calibrate the LLM using a small set of human judgements. Our method ensures a more robust and human-aligned scoring of counterspeech using LLMs.
@@ 3 Dataset
We build and release CSEval, a benchmark dataset for reference-free and multi-dimensional counterspeech evaluation. CSEval contains expert human assessments of $7,926$ model-generated CS, across four quality dimensions (or aspects). In this section, we give a detailed overview of the data curation process, evaluation dimensions, and the models used in CSEval.
@@ 3.3 Models
We include counterspeeches generated using both supervised and prompt-based methods. In terms of supervised methods, we include three popular counterspeech generation models – QUARC Gupta et al. (2023), Zhu and Bhat (2021), DialoGPT Zhang et al. (2020b), and Generate-Prune-Select (GPS). As our prompting baselines, we include counterspeeches generated by two popular LLMs –GPT-3.5-Turbo (ChatGPT) and GPT-4 Ouyang et al. (2022). For each of them,we include both zero- and few-shot prompting baselines. Further details surrounding model training and prompting are provided under Appendix A.1.1.
@@ 3.4 Annotation Process
We used Krippendorff’s alpha coefficient Klaus (2011) to measure the inter-annotator agreement of the expert annotations. For the first round of annotations, we obtained a decent inter-annotator interval of Krippendorff’s (averaged across the four dimensions) of $0.473$. However, the 2nd round increases the overall inter-annotator agreement with an average Krippendoff coefficient of $0.681$. We also calculate the standard deviation of annotator scores within the respective groups. We plot the histogram of these statistics in Appendix Figure 3. Here, we observe that suitableness remains the most contentious dimension among experts. Further details about the annotation process are provided under Appendix Section A.
Figure 2: An overview of the multi-phase auto-calibration framework of Auto-CSEval, including the generation and refinement of CoT instructions, evaluation criteria formulation, and the iterative calibration process aligned with expert human judgement.
@@ 4 Proposed Method
Auto Evaluation Steps (CoT): CoT is a prompt engineering technique that encourages LLMs to break down a complex task into a sequence of intermediate steps and has been shown to improve LLMs in their reasoning abilities and generate more accurate and informative responses for multi-step problems, such as arithmetic, common-sense, and symbolic reasoning Wang et al. (2023); Zheng et al. (2023). For evaluation tasks, some candidate CoTs need more detailed instructions (evaluation steps). It is generally time-consuming to design such evaluation steps for each task manually. Liu et al. (2023a) found that an LLM can generate such evaluation steps by itself and proposed an auto-CoT approach to evaluate NLG tasks. However, this approach may result in sub-optimal and misaligned CoT where the scoring guidelines are absent and only output categorical ranges (e.g., 0-5) are provided. LLM-based evaluators are shown to suffer from insufficient prompting, resulting in inconsistent and misaligned evaluations Lu et al. (2023). However, manually defining CoT instructions or evaluation steps is both challenging and time-consuming and may require an expert who has the requisite domain knowledge to establish the evaluation criteria. Furthermore, manually defining an evaluation criteria may also lead to personal bias of the evaluator. Recently, Liu et al. (2023b) proposed a framework to auto-calibrate LLM prompts based on a validation set of human judgments. Drawing inspiration from this, we propose a data-driven approach to calibrating the most optimal CoT instructions (or evaluation steps) for a given dimension. As illustrated in Figure 2, our method follows a two-phase procedure.
In the first phase, we begin by constructing a validation set $D^{*}$ of $500$ model outputs by randomly sampling from the CSEval dev set. This is denoted by $D$. After this step, we prompt an LLM to independently generate the scoring CoT $C$ from few-shot HS-CS exemplars. While doing this, it is also important to make sure that the generated CoT are non-repetitive. Monte-Carlo Hastings (1970) trials are shown to help in countering any kind of label- or position-related bias during sampling. Therefore, following Liu et al. (2023b), we conduct Monte-Carlo trials and sample fewshot exemplars from the development set $D^{*}$, which are then used directly in the prompt.
Given an arbitrary sample $d_{i}\in D$, the prompt template for chain of thought (CoT) drafting $T_{D}$, evaluation dimension $a$ (e.g., relevance, aggressiveness), and a counterspeech-model $LLM(\cdot)$, the quality dimension $d_{i}$ is evaluated as $\hat{s}_{i,a}=LLM(T_{D}(d_{i},C,a))$. With this setup in place, a corresponding candidate CoT can be inferred as follows:
@@
In the given equation, $D_{s}=\bigcup_{i}(d_{i}^{*},s_{i,a})\subset D^{*}$ represents the subset of few-shot exemplars. On the other hand, $P_{\theta}$ denotes the probability distribution modeled by the LLM’s parameters $\theta$. To allow for diversity in candidate CoT prompts, we apply temperature sampling to draw scoring CoT from the LLM, similar to Liu et al. (2023b). The temperature values for temperature sampling are dynamically selected from a range of (00 to $0.5$), for generation diversity while curating the candidate CoT. The prompt templates used in this process are provided in Appendix B. After obtaining the initial set of candidate CoTs in the first phase, we move to evaluation and refinement in the second phase.
In the second phase, we iteratively refine the candidate CoT pool obtained from the previous phase using $D^{*}$. The CoTs obtained in the first phase may be sub-optimal. In order to filter out high-quality candidates, we first evaluate them using the dev set $D^{*}$. We then sample only the best-performing candidate CoT $C$ based on their Spearman correlation scores (the higher the better). To address the low correlation between human ratings and the initial CoT, we prompt the LLM to refine its previously generated CoT through self-editing. In this process, we use examples with the lowest correlation scores (strong disagreement). Specifically, we identify the top-3 instances with the highest absolute difference between the predicted and human scores and use these as fewshot examples for the CoT-refinement step.. This self-editing step ensures that all candidate CoT $C$ are progressively aligned with human ratings at each iteration. This improves diversity and helps to reduce any biases in the final prompt Liu et al. (2023b).
@@
    Modify: Alter parts of the generated chain-of-thought evaluation steps (CoT) with the aim of increasing its correlation with human scores.
@@
    Paraphrase: If a candidate chain-of-thought (CoT) scores below the correlation threshold, paraphrase it to make it clearer and more concise.
@@
    Addition of new rules or evaluation steps: If the model $LLM(T_{D}(d_{i},C,a))$ finds new scoring rules that are not present in the current candidate CoT $C$, append $C$ with the new rules.
@@
Figure 2 provides an overview of our two-phased calibration process. As shown in the figure, after obtaining a new candidate CoT $C$, we sample them using the validation set $D^{*}$. After this step, we combine the new CoT $\hat{C}$ with the pre-filtered draft CoTs. This gives us a final calibrated set of scoring rules (evaluation steps). Algorithm 1 details the pseudo-code followed for CoT selection and iterative refinement.
@@
GPT-4 (Auto-CSEval) $0.687$ $0.637$ $0.447$ $0.425$ $0.600$ $0.506$ $0.567$ $0.504$ $0.575$ $0.518$
$\Delta_{\mathrm{GPT-4(Auto-CSEval)(Ours)}-\mathrm{BestMethod}}$ $\uparrow 0.062$ $\uparrow 0.065$ $\uparrow 0.068$ $\uparrow 0.075$ $\uparrow 0.015$ $\uparrow 0.019$ $\uparrow 0.018$ $\uparrow 0.040$ $\uparrow 0.048$ $\uparrow 0.049$
Table 1: Sample-level Spearman ($\rho$) and Kendall-Tau ($\tau$) correlations of different metrics on CSEval benchmark. For any given evaluation aspect (Relevance, for example), the correlation scores are computed against their respective human ratings. LLM-based evaluators are reported in three settings: zero-shot, chain-of-thoughts (G-EVAL), and auto-calibrated chain-of-thoughts (Auto-CSEval). Automatic and human-evaluation scores were normalised to a scale of (0,1) before computing correlation.